OpenStudy (anonymous):

find the missing side length. tell if the side lengths form a Pythagorean triple. explain.

OpenStudy (anonymous):

missing side length: 32.01

OpenStudy (anonymous):

@sdfgsdfgs

Nnesha (nnesha):

\[\huge\rm a^2 + b^2 = c^2\] c is longer side of the right triangle

OpenStudy (anonymous):

so it does form a Pythagorean triple?

OpenStudy (sdfgsdfgs):

correct n sorry 32.01 is not the right ans...

OpenStudy (anonymous):

but 25^2 + 20^2 = c^2 625 + 400 = c^2 sqrt1025 = sqrt c^2 32.01=c

Nnesha (nnesha):

hmm i said longer side of right triangle is C hypotenuse

OpenStudy (acxbox22):